Lots on the go.
Rigging of the control surfaces was 100% complete... then realized that the balls on the ends of the flap control cables had come out of their sockets. Fixed. Still need to check tension and safety those cables, but all of the rest are complete (ailerons, elevator, rudder).
Trim tabs set, calibrated on the EFIS and deflection measured.
Shims for main gear toe ordered and received... need longer bolts. Should arrive Feb 3.
Magnetometer / compass swing: My Dynon magnetometer was calibrated backwards - sent for recalibration. Hopefully it will be back this week.
Cleaned fuel filters.
Fuel tank calibration partially done - right wing is complete. May need to redo as sensors are sticky it seems.
Installing the autopilot bridles today. Then that can be calibrated.
Placards started.
Fuel control system checks end to end for jam nuts, cotter pins, etc underway.